Can someone please help me understand graph's before my final next week.
I have 2 questions.
1) It shows a graph given by a quadratic function of the form
f(x)=ax^2, which of these 3 functions matches. The graph shows
a staight line going thru Y at (-2,2) & (2,2)
A)f(x)=2x^2
B)f(x)=3x^2
C)f(x)=1/2x^2
____________________________________
2) It shows a graph is froma quadratic function of the form
f(x)=(x-h)^2 with a symmetry x=-3 and vertex (-3,0). Which
of these 3 matches.
A)f(x)=(x-3)^2
B)f(x)=(x+3)^2
C)f(x)=(x+2)^2
Thanks for all your help

Found 2 solutions by rapaljer, Nate:
Answer by rapaljer(4671)

You don't have to draw a graph, although if you have one in front of you in the problem, it will help. Just take the given values of x, substitute these into the three given functions, and see which function gives you the correct y values.
1. If x = -2, then y = 2, and if x = 2, then y = 2.

A) f(x) = 2x^2
-> f(-2) = 2(-2)^2 = 8. This does NOT work, so it is eliminated.

B) f(x) = 3x^2
-> f(-2) = 3(-2)^2 = 12. Again, this does NOT work.

C) f(x) = 1/2 x^2
-> f(-2) = 1/2* (-2)^2 = 1/2*4= 2. This one checks.
-> f(2) = 1/2* (2)^2 = 1/2*4= 2. This one also checks.

Therefore the answer is C.

Likewise you can do the second problem, right???

R^2 at SCC


Answer by Nate(3500)

The second is just a lot like the first.
f(x)=(x+3)^2
f(-3)=(-3+3)^2
f(-3)=0
(B) is your second answer.